theronhoward400:

Graph the system of inequalities presented here on your own paper, then use your graph to answer the following questions: y > −4x − 1 y is less than 3 over 2 times x minus 1 Part A: Describe the graph of the system, including shading and the types of lines graphed. Provide a description of the solution area. (6 points) Part B: Is the point (−1, −1) included in the solution area for the system? Justify your answer mathematically. (4 points) (10 points)

Extrinix:

Let's firt convert \(y~is~less~than~3~over~2~times~x~minus~1\) \(y< \dfrac{3}{2} x-1\) Ok so, when you graph -4x-1 (y=-4x-1) it shows the right side of the line as shaded, and when you graph \(y< \dfrac{3}{2} x-1\) it shows the right side as well. To figure out Answer A, give a description of what the graph looks like, how the line looks, if it's positive or negative, which sides of it are shaded, etc. To figure out Answer B, take a look at the graph of both of the equations, and look to see if \( (-1,-1) \) exists inside of the shaded areas, not the open areas.